OpenStudy (anonymous):

hat is the solution to the equation? c + (–2) = –0.7 c = 0.6

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. c−2=−0.7 Step 1: Add 2 to both sides. c−2+2=−0.7+2 c=1.3 Answer: c=1.3

OpenStudy (anonymous):

so its 1.3

OpenStudy (ryanwhite499):

Yes it is @fallengoddess